OpenStudy (anonymous):

Which expression can be used to find the area of the circle below? its 12m

OpenStudy (anonymous):

a. A = (3.14)(62) b. A = (3.14)(122) c. A = (3.14)(2 • 6) d. A = (3.14)(2 • 12)

OpenStudy (akashdeepdeb):

You remember the formula I just posted a minute ago in the previous question? Try using that here. :)

OpenStudy (anonymous):

what is 12 m , the radius or the diameter?

OpenStudy (akashdeepdeb):

Area of a circle: \(\pi * r^2\) r = radius. PS Getting the right answer isn't going to help you as much as understanding the concept is.

OpenStudy (anonymous):

not sure if the blank space in between means (2 12 ) means 12 to the power of 2